# Masked Generative Priors Improve World Models Sequence Modelling Capabilities

### v3

@@ Abstract
Deep Reinforcement Learning (RL) has become the leading approach for creating artificial agents in complex environments. Model-based approaches, which are RL methods with world models that predict environment dynamics, are among the most promising directions for improving data efficiency, forming a critical step toward bridging the gap between research and real-world deployment. In particular, world models enhance sample efficiency by learning in imagination, which involves training a generative sequence model of the environment in a self-supervised manner. Recently, Masked Generative Modelling has emerged as a more efficient and superior inductive bias for modelling and generating token sequences. Building on the Efficient Stochastic Transformer-based World Models (STORM) architecture, we replace the traditional MLP prior with a Masked Generative Prior (e.g., MaskGIT Prior) and introduce GIT-STORM. We evaluate our model on two downstream tasks: reinforcement learning and video prediction. GIT-STORM demonstrates substantial performance gains in RL tasks on the Atari 100k benchmark. Moreover, we apply Transformer-based World Models to continuous action environments for the first time, addressing a significant gap in prior research. To achieve this, we employ a state mixer function that integrates latent state representations with actions, enabling our model to handle continuous control tasks. We validate this approach through qualitative and quantitative analyses on the DeepMind Control Suite, showcasing the effectiveness of Transformer-based World Models in this new domain. Our results highlight the versatility and efficacy of the MaskGIT dynamics prior, paving the way for more accurate world models and effective RL policies.
@@ 1 Introduction
Deep Reinforcement Learning (RL) has emerged as the premier method for developing agents capable of navigating complex environments. Deep RL algorithms have demonstrated remarkable performance across a diverse range of games, including arcade games (Mnih et al., 2015; Schrittwieser et al., 2020; Hafner et al., 2021; 2023), real-time strategy games (Vinyals et al., 2019; OpenAI, 2018), board games (Silver et al., 2016; 2018; Schrittwieser et al., 2020), and games with imperfect information (Schmid et al., 2021). Despite these successes, data efficiency remains a significant challenge, impeding the transition of deep RL agents from research to practical applications. Accelerating agent-environment interactions can mitigate this issue to some extent, but it is often impractical for real-world scenarios. Therefore, enhancing sample efficiency is essential to bridge this gap and enable the deployment of RL agents in real-world applications (Micheli et al., 2022). 1 Delft University of Technology, NL. 2 Bangladesh University of Engineering and Technology, BD. 3 The University of Tokyo, JP. 4 Mila, University of Montreal, CA. 5 National University of Singapore, SG. Corresponding author: c.meo@tudelft.nl
@@ 2.1 Model-based RL: World Models
With the advent of transformers (Vaswani et al., 2017) in sequence modelling and the promise of scaling performance across multiple tasks with more data, replacing the traditional RSSM backbones with transformer-based backbones has become a very active research direction. IRIS (Micheli et al., 2022) was one of the first transformer-based world model approaches, using VQVAE (Oord et al., 2017) to tokenize the observations and an autoregressive transformer that performs sequence modelling in the token space. Although IRIS obtains human-level performance on 10 out of 26 Atari games, its actor-critic operates in the RGB pixel space, making it almost 14x slower than DreamerV3. In contrast, methods such as TWM (Robine et al., 2023) and STORM (Zhang et al., 2023), use latent actor-critic input space. The proposed GIT-STORM employs it as well, as we believe it is the most promising direction to overcome sample efficiency constraints. TWM introduces a balanced sampling scheme and employs TransformerXL (Dai et al., 2019). More recently, STORM updated DreamerV3 by utilizing the transformer backbone. All aforementioned transformer-based world models use an MLP head to model a dynamics prior which is used to predict the discrete representation of the following timestep. In contrast, introduced by TECO (Yan et al., 2023), we employ a MaskGIT (Chang et al., 2022) prior head, which enhances the sequence modelling capabilities of the world model. Table 2 compares various design aspects of different world models. Furthermore, besides STORM, all the mentioned transformer-based world models concatenate the discrete action to the extracted categorical latent representations. As a result, none of these methods is able to handle continuous actions. In contrast, combining latent representations and actions with a state mixer, we successfully train STORM and GIT-STORM on a challenging continuous action environment (i.e., DMC).
Module DreamerV3 IRIS TWM STORM GIT-STORM
Latent space [OneHot, Hidden] VQ Codes OneHot OneHot OneHot
Dynamics Prior MLP MLP MLP MLP MaskGIT
Experience Sampling Uniform Uniform Balanced Uniform Uniform
@@ 3.2 Dynamics Prior Head: MaskGIT Prior
First, the hidden states, $h_{t}$, are concatenated with the masked latent representations, $z_{t}\circ m_{t}$, where $\circ$ indicates element-wise multiplication. Despite $h_{t}$ being indexed by $t$, it represents the output of the $f_{\theta}$ and thus encapsulates information about the subsequent timestep. Consequently, the concatenation of $z_{t}$ and $h_{t}$ integrates information from both the current and the next timestep, respectively. A bidirectional transformer is then used to learn the relationships between these two consecutive representations, producing a summary representation, $\xi_{t}$. Finally, logits are computed as the dot product (denoted as $\odot$ in Figure 1) between the MaskGIT embeddings, which represent the masked tokens, and $\xi_{t}$. This dot product can be interpreted as a similarity distance between the embeddings and $\xi_{t}$. Indeed, from a geometric perspective, both cosine similarity and the dot product serve as similarity metrics, with cosine similarity focusing on the angle between two vectors, while the dot product accounts for both the angle and the magnitude of the vectors. Therefore, by optimizing the MaskGIT prior, this dot product aligns the embeddings with $\xi_{t}$, thereby facilitating and improving the computation of logits. In contrast, when using the MLP prior, the logits are generated as the output of an MLP that only takes $h_{t}$ as input. This approach requires the model to learn the logits space and their underlying meaning without any inductive bias, making the learning process more challenging.
@@ 3.3 State Mixer for Continuous Action Environments
Algorithm 1 Draft-and-Revise decoding scheme
2: /* draft phase */
[…]
4:$\bm{\Pi}\sim p_{\text{draft}}(\bm{\Pi};T_{\text{draft}})$
5: /* generate a draft prior map */
6:$\bm{z}^{0}\leftarrow\textsc{Update}(\bm{z}^{\text{empty}},\bm{\Pi},h_{t};\theta)$
7: /* revision phase */
8:for $\gamma=1,\cdots,\Gamma$ do
9:     $\mathbf{\Pi}\sim p_{\text{revise}}(\mathbf{\Pi};T_{\text{revise}})$
10:     $\bm{z}^{\gamma}\leftarrow\textsc{UPDATE}(\bm{z}^{\gamma-1},\mathbf{\Pi},h_{t};\theta)$
11:end for
12:return $\bm{z}^{\Gamma}$
Algorithm 2 UPDATE of $\bm{z}$
1:Generated latents $\bm{z}$, hidden states $h_{t}$, partition $\mathbf{\Pi}=(\mathbf{m}^{1},\cdots,\mathbf{m}^{T})$, model $\theta$
2: /* update the codes */
3:for $i=1,\cdots,T$ do
4:     $\bm{\hat{z}}\sim p_{\theta}(\bm{\hat{z}}|\bm{z}\circ\mathbf{m}^{i},h_{t})$
5:     $\bm{z}=(1-\mathbf{m}^{i})\circ\bm{z}+\mathbf{m}^{i}\circ\bm{\hat{z}}$
6:end for
7:return $\bm{z}$
@@ 3.4 Imagination Phase
Differently from STORM, which uses a MLP prior to compute the next timestep representations, we employ MaskGIT to accurately model the dynamics of the environment. By improving the quality of the predicted trajectories, the agent is able to learn a superior policy. Moreover, while STORM applies KV caching only on the conditioning phases, we extend its use to the whole imagination phase as in IRIS (Micheli et al., 2022), significantly improving the related sample efficiency.
@@ 4.1 Experimental Setup
Atari 100k benchmark consists of 26 different video games with discrete action space. The constraint of 100k interactions corresponds to a total of 400k frames used for training, as frame skipping is set to 4. For RL task on Atari 100k benchmark, we compare GIT-STORM against one model-free method, SimPLe (Kaiser et al., 2019), one RSSM, DreamerV3 (Hafner et al., 2023), and three TSSM models (i.e., IRIS (Micheli et al., 2022), TWM (Robine et al., 2023), and STORM (Zhang et al., 2023)). DMC benchmark consists of 18 control tasks with continuous action space. We restrict the models to be trained with only 500k interactions (1M frames) by setting frame skipping to 2. For RL task on DMC benchmark, we compare our model against SAC (Haarnoja et al., 2018), CURL (Laskin et al., 2020), DrQ-v2 (Yarats et al., 2022), PPO (Schulman et al., 2017), DreamerV3 (Hafner et al., 2023), and STORM (Zhang et al., 2023). We trained GIT-STORM on 5 different seeds. For video prediction tasks, we compare GIT-STORM with STORM only to understand how the MaskGIT Prior affects the visual quality of predicted frames and its influence on the policy training.
Extended details of the baselines for both benchmarks can be found in Appendix I.
@@
Proper evaluation of RL algorithms is known to be difficult due to both the stochasticity and computational requirements of the environments  (Agarwal et al., 2021). To provide an accurate evaluation of the models, we consider a series of metrics to assess the performances of the considered baselines on across the selected experiments. We report human normalized mean and median as evaluation metrics, aligning with prior literature. We also report interquartile Mean (IQM), Optimality Gap, Performance Profiles (scores distributions), and Probability of Improvement (PI), which provide a statistically grounded perspective on the model evaluation (Agarwal et al., 2021).
For video prediction task, we report two metrics: Fréchet Video Distance (FVD) (Unterthiner et al., 2019) to evaluate visual quality of the predicted frames, and perplexity (Jelinek et al., 2005) measure of the predicted tokens to evaluate the token utilization by the dynamics prior head. We use the trained agent to collect ground truth episodes and use the world model to predict the frames. We report the FVD over 256 videos which are conditioned on the first 8 frames to predict 48 frames.
A full description of these metrics can be found in Appendix J.
@@ 5 Discussion
The proposed GIT-STORM uses a Masked Generative Prior (MaskGIT) to enhance the world model sequence modelling capabilities. Indeed, as discussed in the introduction, high quality and accurate representations are essential to guarantee and enhance agent policy learning in imagination. Remarkably, the proposed GIT-STORM is the only world model, among the ones that use uniform sampling and latent actor critic input space, that is able to achieve non-zero reward on the Freeway environment (e.g., DreamerV3: 0, STORM: 0 $\rightarrow$ GIT-STORM: 13). Indeed, both STORM and IRIS resorted in ad-hoc solutions to get positive rewards, such as changing the sampling temperature (Micheli et al., 2022) and using demonstration trajectories (Zhang et al., 2023). Such result, together with the quantitative results on the Atari 100k and DMC benchmarks, clearly answer question (a) - the presented MaskGIT prior improves the policy learning behavior and performance on downstream tasks (e.g., Model-based RL and Video Prediction) of TSSMs. Moreover, the FVD and perplexity comparisons in Table 4 and Table 4 suggest that GIT-STORM has better predictive capabilities, learns a better dynamics module, and presents more accurate imagined trajectories (Figure 5, Figure 6). Similarly to image synthesis (Chang et al., 2022) and video prediction (Yu et al., 2023a) tasks, we show how using masked generative modelling is a better inductive bias to model the prior dynamics of discrete representations and improve the downstream usefulness of world models on RL tasks. Furthermore, the MaskGIT Prior can be used in any sequence modelling backbone that uses categorical latent representations (e.g., VideoGPT (Yan et al., 2021), IRIS (Micheli et al., 2022)), positioning itself as a very versatile approach. In this work we do not apply a MaskGIT prior on top of IRIS only because of computational and time constraints - IRIS requires 168h of training on a V100 GPU for a single run (Zhang et al., 2023) .
@@ Appendix C Full Results on RL Task
Figure 10: More generated trajectories of STORM and GIT-STORM on selected DMC environments. The model uses the first 8 frames as context and then generates the following 48 frames.
@@ Appendix D Training Curves
Figure 11: Performance evaluation across all the checkpoints for the Atari 100k benchmark. The solid line represents the average over $5$ seeds while the fill area is defined in terms of maximum and minimum values corresponding to each checkpoint.
@@ Appendix E Ablation study
E.1 Dimensions of Dynamic Prior Head

### v5

@@ Abstract
Deep Reinforcement Learning (RL) has become the leading approach for creating artificial agents in complex environments. Model-based approaches, which are RL methods with world models that predict environment dynamics, are among the most promising directions for improving data efficiency, forming a critical step toward bridging the gap between research and real-world deployment. In particular, world models enhance sample efficiency by learning in imagination, which involves training a generative sequence model of the environment in a self-supervised manner. Recently, Masked Generative Modelling has emerged as a more efficient and superior inductive bias for modelling and generating token sequences. Building on the Efficient Stochastic Transformer-based World Models (STORM) architecture, we replace the traditional MLP prior with a Masked Generative Prior (e.g., MaskGIT Prior) and introduce GIT-STORM. We evaluate our model on two downstream tasks: reinforcement learning and video prediction. GIT-STORM demonstrates substantial performance gains in RL tasks on the Atari 100k benchmark. Moreover, we apply Categorical Transformer-based World Models to continuous action environments for the first time, addressing a significant gap in prior research. To achieve this, we employ a state mixer function that integrates latent state representations with actions, enabling our model to handle continuous control tasks. We validate this approach through qualitative and quantitative analyses on the DeepMind Control Suite, showcasing the effectiveness of Transformer-based World Models in this new domain. Our results highlight the versatility and efficacy of the MaskGIT dynamics prior, paving the way for more accurate world models and effective RL policies. Here you can find our GitHub repository: https://github.com/Cmeo97/GitStorm/.
@@ 1 Introduction
Deep Reinforcement Learning (RL) has emerged as the premier method for developing agents capable of navigating complex environments. Deep RL algorithms have demonstrated remarkable performance across a diverse range of games, including arcade games (Mnih et al., 2015; Schrittwieser et al., 2020; Hafner et al., 2021; 2023), real-time strategy games (Vinyals et al., 2019; OpenAI, 2018), board games (Silver et al., 2016; 2018; Schrittwieser et al., 2020), and games with imperfect information (Schmid et al., 2021). Despite these successes, data efficiency remains a significant challenge, impeding the transition of deep RL agents from research to practical applications. Accelerating agent-environment interactions can mitigate this issue to some extent, but it is often impractical for real-world scenarios. Therefore, enhancing sample efficiency is essential to bridge this gap and enable the deployment of RL agents in real-world applications (Micheli et al., 2022; Lanillos et al., 2021). 1 Delft University of Technology, NL. 2 Bangladesh University of Engineering and Technology, BD. 3 The University of Tokyo, JP. 4 Mila, University of Montreal, CA. 5 National University of Singapore, SG. Corresponding author: c.meo@tudelft.nl
@@ 2.1 Model-based RL: World Models
With the advent of transformers (Vaswani et al., 2017) in sequence modelling and the promise of scaling performance across multiple tasks with more data, replacing the traditional RSSM backbones with transformer-based backbones has become a very active research direction. Although IRIS  (Micheli et al., 2022), one of the first transformer-based world model approaches, obtains impressive results, its actor-critic operates in the RGB pixel space, making it almost 14x slower than DreamerV3. In contrast, methods such as TWM (Robine et al., 2023) and STORM (Zhang et al., 2023), use latent actor-critic input space. The proposed GIT-STORM employs it as well, as we believe it is the most promising direction to overcome sample efficiency constraints. More recently, STORM updated DreamerV3 by utilizing the transformer backbone. All aforementioned transformer-based world models use an MLP head to model a dynamics prior which is used to predict the discrete representation of the following timestep. In contrast, introduced by TECO (Yan et al., 2023), we employ a MaskGIT (Chang et al., 2022) prior head, which enhances the sequence modelling capabilities of the world model. Table 4 compares various design aspects of different world models. Furthermore, besides STORM, all the mentioned transformer-based world models concatenate the discrete action to the extracted categorical latent representations. As a result, none of these methods is able to handle continuous actions. In contrast, combining latent representations and actions with a state mixer, we successfully train STORM and GIT-STORM on a challenging continuous action environment (i.e., DMC).
@@ 3.2 Dynamics Prior Head: MaskGIT Prior
First, the hidden states, $h_{t}$, are concatenated with the masked latent representations, $z_{t}\circ m_{t}$, where $\circ$ indicates element-wise multiplication. Despite $h_{t}$ being indexed by $t$, it represents the output of the $f_{\theta}$ and thus encapsulates information about the subsequent timestep. Consequently, the concatenation of $z_{t}$ and $h_{t}$ integrates information from both the current and the next timestep, respectively. A bidirectional transformer is then used to learn the relationships between these two consecutive representations, producing a summary representation, $\xi_{t}$. Finally, logits are computed as the dot product (denoted as $\odot$ in Figure 1) between the MaskGIT embeddings, which represent the masked tokens, and $\xi_{t}$. This dot product is also known as weight tying strategy, first formalized in Inan et al. (2017) and then used in the original MaskGIT Chang et al. (2022) and GPT-2 Radford et al. (2019) models as well because of its regularization effects that help preventing overfitting Inan et al. (2017). Indeed, this weight tying strategy (i.e., dot product) can be interpreted as a similarity distance between the embeddings and $\xi_{t}$. Indeed, from a geometric perspective, both cosine similarity and the dot product serve as similarity metrics, with cosine similarity focusing on the angle between two vectors, while the dot product accounts for both the angle and the magnitude of the vectors. Therefore, by optimizing the MaskGIT prior, this dot product aligns the embeddings with $\xi_{t}$, thereby facilitating and improving the computation of logits. In contrast, when using the MLP prior, the logits are generated as the output of an MLP that only takes $h_{t}$ as input. This approach requires the model to learn the logits space and their underlying meaning without any inductive bias, making the learning process more challenging.
@@ 3.4 Imagination Phase
Differently from STORM, which uses a MLP prior to compute the next timestep representations, we employ MaskGIT to accurately model the dynamics of the environment. By improving the quality of the predicted trajectories, the agent is able to learn a superior policy.
@@ 4.1 Experimental Setup
Atari 100k benchmark consists of 26 different video games with discrete action space. The constraint of 100k interactions corresponds to a total of 400k frames used for training, as frame skipping is set to 4. For RL task on Atari 100k benchmark, we compare GIT-STORM against one model-free method, SimPLe (Kaiser et al., 2019), one RSSM, DreamerV3 (Hafner et al., 2023), and three TSSM models (i.e., IRIS (Micheli et al., 2022), TWM (Robine et al., 2023), and STORM (Zhang et al., 2023)). DMC benchmark consists of 18 control tasks with continuous action space. We restrict the models to be trained with only 500k interactions (1M frames) by setting frame skipping to 2. For RL task on DMC benchmark, we compare our model against SAC (Haarnoja et al., 2018), CURL (Laskin et al., 2020), DrQ-v2 (Yarats et al., 2022), PPO (Schulman et al., 2017), DreamerV3 (Hafner et al., 2023), and STORM (Zhang et al., 2023). We trained GIT-STORM on 5 different seeds. For video prediction tasks, we compare GIT-STORM with STORM only to understand how the MaskGIT Prior affects the visual quality of predicted frames and its influence on the policy training. Extended details of the baselines for both benchmarks can be found in Appendix J.
@@
Proper evaluation of RL algorithms is known to be difficult due to both the stochasticity and computational requirements of the environments  (Agarwal et al., 2021). To provide an accurate evaluation of the models, we consider a series of metrics to assess the performances of the considered baselines on across the selected experiments. We report human normalized mean and median as evaluation metrics, aligning with prior literature. We also report interquartile Mean (IQM), Optimality Gap, Performance Profiles (scores distributions), and Probability of Improvement (PI), which provide a statistically grounded perspective on the model evaluation (Agarwal et al., 2021).For video prediction task, we report two metrics: Fréchet Video Distance (FVD) (Unterthiner et al., 2019) to evaluate visual quality of the predicted frames, and perplexity (Jelinek et al., 2005) measure of the predicted tokens to evaluate the token utilization by the dynamics prior head. We use the trained agent to collect ground truth episodes and use the world model to predict the frames. We report the FVD over 256 videos which are conditioned on the first 8 frames to predict 48 frames. A full description of these metrics can be found in Appendix K.
@@ 5 Discussion
The proposed GIT-STORM uses a Masked Generative Prior (MaskGIT) to enhance the world model sequence modelling capabilities. Indeed, as discussed in the introduction, high quality and accurate representations are essential to guarantee and enhance agent policy learning in imagination. Remarkably, the proposed GIT-STORM is the only world model, among the ones that use uniform sampling and latent actor critic input space, that is able to achieve non-zero reward on the Freeway environment (e.g., DreamerV3: 0, STORM: 0 $\rightarrow$ GIT-STORM: 13). Indeed, both STORM and IRIS resorted in ad-hoc solutions to get positive rewards, such as changing the sampling temperature (Micheli et al., 2022) and using demonstration trajectories (Zhang et al., 2023). Such result, together with the quantitative results on the Atari 100k and DMC benchmarks, clearly answer question (a) - the presented MaskGIT prior improves the policy learning behavior and performance on downstream tasks (e.g., Model-based RL and Video Prediction) of TSSMs. Moreover, the FVD and perplexity comparisons in Table 3 and Table 3 suggest that GIT-STORM has better predictive capabilities, learns a better dynamics module, and presents more accurate imagined trajectories (Figure 10, Figure 11). Similarly to image synthesis (Chang et al., 2022) and video prediction (Yu et al., 2023a) tasks, we show how using masked generative modelling is a better inductive bias to model the prior dynamics of discrete representations and improve the downstream usefulness of world models on RL tasks. Furthermore, the MaskGIT Prior can be used in any sequence modelling backbone that uses categorical latent representations (e.g., VideoGPT (Yan et al., 2021), IRIS (Micheli et al., 2022)), positioning itself as a very versatile approach. In this work we do not apply a MaskGIT prior on top of IRIS only because of computational and time constraints.
@@ A.2 Policy Learning
A.3 Draft and Revise decoding scheme
Algorithm 1 Draft-and-Revise decoding scheme
2: /* draft phase */
[…]
4:$\bm{\Pi}\sim p_{\text{draft}}(\bm{\Pi};T_{\text{draft}})$
5: /* generate a draft prior map */
6:$\bm{z}^{0}\leftarrow\textsc{MaskGIT Head}(\bm{z}^{\text{empty}},\bm{\Pi},h_{t};\theta)$
7: /* revision phase */
8:for $\gamma=1,\cdots,\Gamma$ do
9:     $\mathbf{\Pi}\sim p_{\text{revise}}(\mathbf{\Pi};T_{\text{revise}})$
10:     $\bm{z}^{\gamma}\leftarrow\textsc{MaskGIT Head}(\bm{z}^{\gamma-1},\mathbf{\Pi},h_{t};\theta)$
11:end for
12:$\bm{z}_{t+1}\leftarrow\bm{z}^{\Gamma}$
13:return $\bm{z}_{t+1}$
Algorithm 2 MaskGIT Head
1:Generated latents $\bm{z}$, hidden states $h_{t}$, partition $\bm{\Pi}=(\mathbf{m}^{1},\ldots,\mathbf{m}^{T})$, model $\theta$
2$\triangleright$ Update the codes
3:for $i=1$ to $T$ do
[…]
5:     $\xi\leftarrow\operatorname{BidirectionalTransformer}(\text{MaskGIT\_codes},h_{t})$
[…]
7:     $\hat{\bm{z}}\sim\operatorname{Categorical}(\text{logits})$
8:     $\bm{z}\leftarrow(1-\mathbf{m}^{i})\circ\bm{z}+\mathbf{m}^{i}\circ\hat{\bm{z}}$
9:end for
10:return $\bm{z}$
@@ Appendix B Extended Related Works: Video Prediction Modelling
Module DreamerV3 IRIS TWM STORM GIT-STORM
Latent space [OneHot, Hidden] VQ Codes OneHot OneHot OneHot
Dynamics Prior MLP MLP MLP MLP MaskGIT
Experience Sampling Uniform Uniform Balanced Uniform Uniform
@@ Appendix D Training Curves
Figure 8: Training profiles across all the checkpoints for the Atari 100k benchmark. The solid line represents the average over $5$ seeds while the fill area is defined in terms of maximum and minimum values corresponding to each checkpoint.
@@ Appendix E Ablation study
E.1 GIT-STORM Ablations
In this section, we analyze the contributions of the two primary components that define GIT-STORM:
  • MaskGIT Head: We compare the performance of the MaskGIT head against a standard MLP head to assess its role in improving downstream results.
  • Logits Computation via Dot Product: We evaluate the impact of computing logits as the dot product between $\xi_{t}$ and the MaskGIT embeddings, comparing this approach to the alternative of using an MLP head that takes $\xi_{t}$ as input and directly outputs logits.
These components are hypothesized to be critical for understanding the capabilities of GIT-STORM and the individual contributions they make to the observed performance improvements.
Figure 12 illustrates an ablation study on three Atari games (Hero, Freeway, and Boxing) and three DMC environments (Walker Walk, Walker Run, and Quadruped Run). Across both sets of environments, the removal of the MaskGIT head consistently results in poorer downstream performance (e.g., lower scores). Additionally, leveraging the dot product between $\xi_{t}$ and MaskGIT embeddings has a substantial impact in environments such as Freeway, Walker Walk, and Quadruped Run. However, its influence appears negligible in other environments like Hero and Walker Run, suggesting that its efficacy may be context-dependent.
Figure 12: GIT-STORM ablation study on selected Atari and DMC environments: GIT-STORM w/o MG stands for without MaskGIT head, while GIT-STORM w/o MGDP stands for without MaskGIT dot product. All results are averaged across three random seeds.
E.2 Dimensions of Dynamic Prior Head